How fast would you be driving if you went 3 miles in 19 minutes?
Thank you very much,
Marlene 9/18/36

Answer by rosey51235(30)

Hi :)
If by how fast you mean in Miles Per Hour, then you would set this up in a proportion.
(/=fraction bar)
3/19=x/60 (3 and x correspond because that is a miles measurement, while 19 and 60 correspond because they are a minutes measurement)
Cross multiply---198x=19x and 3*60=180
19x=180 --- T oget x alone, divide both sides by 19 (19x/19 will cancel the 19's out, leaving x alone)
180/19=9.4737, though I assume you'd round.
So, 3 miles in 19 minutes is 9 MPH.
....That's reeeeeally slow. even the signs for "keep kids alive"is faster then that, it's "keep kids alive, drive 25"! So yes, 3 miles in 19 minutes=9MPH=very slow.
Hope I helped! :)
-Rose